OpenStudy (anonymous):

I have one last problem :) And then in done!! :) please help!! :D I will post the problem in a comment! :)

OpenStudy (anonymous):

Ok the options are: A. x = 1 B. x = 0.5 C. x = 2

OpenStudy (anonymous):

plessss help? :)

OpenStudy (anonymous):

B X+2 + 3(3-x) = 10 x+2+9-3x=10 -2x=10-11 x=1/2

OpenStudy (anonymous):

Ok so then it would be B thanks a bunch :)